### v1

# AttackSeqBench: Benchmarking Large Language Models’ Understanding of Sequential Patterns in Cyber Attacks

Abstract.
The observations documented in Cyber Threat Intelligence (CTI) reports play a critical role in describing adversarial behaviors, providing valuable insights for security practitioners to respond to evolving threats. Recent advancements of Large Language Models (LLMs) have demonstrated significant potential in various cybersecurity applications, including CTI report understanding and attack knowledge graph construction. While previous works have proposed benchmarks that focus on the CTI extraction ability of LLMs, the sequential characteristic of adversarial behaviors within CTI reports remains largely unexplored, which holds considerable significance in developing a comprehensive understanding of how adversaries operate. To address this gap, we introduce AttackSeqBench, a benchmark tailored to systematically evaluate LLMs’ capability to understand and reason attack sequences in CTI reports. Our benchmark encompasses three distinct Question Answering (QA) tasks, each task focuses on the varying granularity in adversarial behavior. To alleviate the laborious effort of QA construction, we carefully design an automated dataset construction pipeline to create scalable and well-formulated QA datasets based on real-world CTI reports. To ensure the quality of our dataset, we adopt a hybrid approach of combining human evaluation and systematic evaluation metrics. We conduct extensive experiments and analysis with both fast-thinking and slow-thinking LLMs, while highlighting their strengths and limitations in analyzing the sequential patterns in cyber attacks. The overarching goal of this work is to provide a benchmark that advances LLM-driven CTI report understanding and fosters its application in real-world cybersecurity operations. Our dataset and code are available at https://github.com/Javiery3889/AttackSeqBench.
Cyber Security, Cyber Threat Intelligence, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacyccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning approaches
1. Introduction
Figure 1. Illustration of the difference between existing CRU-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl and Hannon, 2020). We highlight the sequential patterns within CTI reports and investigate the capability of LLMs in comprehending and reasoning such sequential patterns.
The increasing complexity and diversity of cyber attacks have imposed a challenge to defend against Advanced Persistent Threats (APTs) (Huang et al., 2018; Duo et al., 2022). Cyber Threat Intelligence (CTI) reports, which detail the adversarial behaviors observed in cyber attacks, are crucial towards addressing emerging cyber threats. However, their unstructured nature demands extensive manual analysis (Wagner et al., 2019; Sun et al., 2023). Meanwhile, recent advancements in Large Language Models (LLMs) have demonstrated promising applications in cybersecurity, such as attack knowledge graph construction (Zhang et al., 2025b; Hu et al., 2024) and threat intelligence extraction (Rahman et al., 2023; Xu et al., 2024b). These capabilities shed new light on LLM-driven CTI Report Understanding (CRU) to support cybersecurity operations, leading to the development of various CRU-related benchmarks to assess LLMs’ capability in the field.
Existing CRU-related benchmarks primarily focus on evaluating LLMs on three aspects: (1) Classifying adversarial behaviors (Alam et al., 2023) (2) Extracting threat intelligence from unstructured text (Bhusal et al., 2024), or (3) Understanding meaningful correlations between the cyber threats and malware groups (Alam et al., 2024). While these benchmarks focus on the information extraction capabilities of LLMs, their ability to understand and reason about the sequential patterns of adversarial behavior remains largely unexplored. We refer to these sequences as attack sequences (Choi et al., 2021; Al-Sada et al., 2025), which represent the temporal order of malicious actions adversaries perform to achieve their goals. The actions which constitute an attack sequence are commonly referred to as attack patterns, characterizing adversarial behavior on a tactical level (Mavroeidis and Bromander, 2021; Siracusano et al., 2023). We illustrate the key differences between existing CRU-related tasks and analyzing attack sequences in Figure 1. Although knowledge bases (KBs) such as the MITRE ATT&CK® framework (Strom et al., 2018) document real-world adversary behaviors through defined attack patterns, analyzing the patterns individually is insufficient to fully capture the progression of cyber attacks as documented in CTI reports. The sophisticated and stealthy nature of APTs requires a comprehensive understanding of how adversaries transition between the different attack phases, which are orchestrated as an attack sequence. This raises the need to consider the sequential characteristics of a cyber attack during CRU.
To address this research gap, we delve into the suitability of LLMs in analyzing attack sequences. We highlight three key perspectives when formulating our main research task:
Perspective 1: The process of mining threat intelligence from CTI reports not only requires general text comprehension; it also necessitates expert-level cybersecurity knowledge. This requirement becomes apparent in attack sequence analysis, as it requires a deep understanding of adversarial behavior to formulate the associate attack sequence. Although LLMs demonstrate significant potential in CRU, prior research has identified limitations in their factual reliability, particularly in technical and knowledge-intensive tasks (Wang et al., 2023a; Xu et al., 2024a). In real-world cybersecurity operations, these limitations can hinder precise and timely analysis of CTI reports, ultimately impacting defensive strategies.
Perspective 2: LLMs have demonstrated remarkable capabilities across a broad range of scenarios, from traditional Natural Language Processing (NLP) tasks to more even more challenging real-world problems (Zhao et al., 2023). The recent breakthrough of slow-thinking reasoning LLMs has also shed new light on complex reasoning tasks related to coding and mathematical reasoning (DeepSeek-AI, 2025; OpenAI, 2025). Yet, the extent to which LLMs can be applied to attack sequence analysis has remained under-explored.
Perspective 3: To address the constantly evolving cyber threat landscape and the advancements of LLMs, the benchmark must be extensible to incorporate newly observed attack sequences from CTI reports. Additionally, the unstructured nature of CTI reports necessitates the benchmark to be generalizable to ensure a comprehensive and realistic evaluation of LLMs’ capabilities. Most importantly, the benchmark must be reproducible to facilitate fair and consistent evaluations across the different LLMs.
Building on these perspectives, we introduce AttackSeqBench, a novel benchmark designed for extensive investigation. Specifically, considering Perspective 1, we define three distinct Question Answering (QA) benchmarking tasks that follow the three-level structure of adversary behaviors —Tactics, Techniques, and Procedures (TTPs) —as outlined in the ATT&CK KB. This hierarchical representation allows security practitioners to represent observed attack patterns in varying levels of granularity based on pre-defined adversarial behaviors. Each tactic represents a high-level objective within a cyber attack’s lifecycle, while techniques and procedures elaborate on the specific actions performed by an adversary to achieve their objectives. Following Perspective 2, we design three benchmark settings to extensively evaluate the effectiveness and robustness of AttackSeqBench. Moreover, we leverage a diverse set of fast-thinking and slow-thinking LLMs in our experiments (Min et al., 2024), providing comprehensive insights into the capabilities and limitations of LLMs in attack sequence analysis. Catering to Perspective 3, we construct attack sequences based on real-world CTI reports. Additionally, we carefully design an automated dataset generation pipeline to generate questions in our QA dataset based on the attack sequence constructed. We summarize our contributions as follow:
  • We propose AttackSeqBench, a novel benchmark designed to systematically evaluate LLMs’ capabilities in analyzing attack sequences constructed from real-world CTI reports.
  • We introduce three QA tasks in our benchmark, each aligned with the different levels of granularity in adversarial behavior representation.
  • We conduct extensive experiments and various benchmarking settings using a diverse set of fast-thinking and slow-thinking LLMs to provide insights into their capabilities and limitations.
2. Related Work
CTI Report Understanding and LLMs. CTI refers to the collection and analysis of evidence-based information related to cyber attacks and threat actors, enabling security practitioners to gain actionable insights into the adversarial operations respond to these adversaries effectively (Wagner et al., 2019; Sun et al., 2023). A crucial source of CTI comes from CTI reports published by security vendors, these reports are typically presented as unstructured textual documents and document observations based on emerging and ongoing cyber threats (Rahman et al., 2020). CTI Report Understanding (CRU) broadly encompasses the extraction, interpretation, and reasoning of key insights within CTI reports. TTP extraction is a critical aspect in CRU, enabling security practitioners to characterize adversarial behaviors into known attack patterns defined in the ATT&CK KB (Husari et al., 2017). Some works (Legoy et al., 2020; Marchiori et al., 2023) leverage multi-label text classification to extract TTPs, while more recent research utilize Language Models (LMs) to perform TTP extraction (Rani et al., 2024). Another line of work (Shin et al., 2023) explores embedding TTPs using Global Vectors for Word Representation to capture co-occurrences. Despite considerable progress, these works tend to be costly and error-prone as adversarial behavior continues evolving.
With the breakthrough of LLMs, researchers have progressively explored their applicability within the cybersecurity domain (Zhang et al., 2024a). Within CRU, existing research (Fayyazi et al., 2024) has showcased significant capabilities of LLMs in interpreting TTPs from the ATT&CK KB, surpassing the performance of some of the fine-tuned LMs with cybersecurity data. Another line of works (Huang and Xiao, 2024; Cheng et al., 2024) utilizes LLMs to extract the relevant entities to construct knowledge graphs (KGs) based on the CTI reports. AttacKG+ (Zhang et al., 2025b) constructs adversarial behavior KGs by leveraging LLMs to map extracted entities to ATT&CK TTPs. However, the capabilities of LLMs to analyze attack sequences in CTI reports remain unexplored.
Related Benchmarks. Within the cybersecurity domain, existing work (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023) have proposed several benchmarks to evaluate LLMs’ cybersecurity knowledge. However, these works primarily rely on authoritative sources (e.g., textbooks, open standards, and example TTPs from ATT&CK KB) and do not incorporate CTI reports in their dataset. CTIBench (Alam et al., 2024) designs four CTI-related tasks, including threat actor attribution and vulnerability assessment, yet only the threat actor attribution task leverages CTI reports. CTISum (Peng et al., 2024) evaluates the ability of LLMs to summarize key facts and attack processes from CTI reports. Extending this work, SevenLLM (Ji et al., 2024) designs an additional task that focuses on extracting entities and relations. To add on, prior research (Siracusano et al., 2023) assesses LLMs’ capabilities in extracting attack patterns, along with entities and relations. In the general domain, existing works (Wang et al., 2023b; Valmeekam et al., 2023) have proposed benchmarks to evaluate the reasoning capabilities of LLMs in sequence-related tasks. Furthermore, previous research has highlighted LLMs capabilities in temporal event analysis (Zhang et al., 2024b; Chang et al., 2024). However, these works may not accurately reflect the extent to which LLMs can reason about attack sequences. In contrast, AttackSeqBench delves into the suitability of LLMs in attack sequence analysis through three tasks that requires in-depth knowledge of adversarial behavior, while emphasizing the practical aspects of CRU by utilizing real-world CTI reports to construct attack sequences in our dataset.
3. Task Definition and Dataset
3.1. Task Definition
An attack sequence $S$ models the progression of adversarial behavior in a cyber attack, characterized by the logical order of the ATT&CK tactics observed in a CTI report alongside the associated techniques and procedures. Formally, $S$ is defined as a 4-tuple:
$S=(T,E,P,O),$
where:
  1. (1)
    Tactic Sequence $T=(t_{1},\ldots,t_{n})$: An ordered list of ATT&CK tactics, which outlines the phases of a cyber attack identified in a CTI report. Note that tactic $t_{k}$ represents the $k$-th tactic.
  2. (2)
    Technique Mapping $E(t_{k})=\{e_{1,k},\ldots,e_{i_{k},k}\}$: A set of ATT&CK techniques (and/or sub-techniques) associated with tactic $t_{k}$. The overall set of techniques across all tactics in the CTI report is represented as $E=\bigcup_{k=1}^{n}E(t_{k})$.
  3. (3)
    Procedure Mapping $P(e_{j,k})=\{p_{1,j,k},\ldots,p_{m_{j,k},j,k}\}$: A set of procedures describing how technique $e_{j,k}$ is performed, the overall set of procedures across all techniques in the CTI report is represented as $P=\bigcup_{e\in E}P(e)$. Each procedure $p_{l,j,k}$ is expressed as a (subject, action, object) triplet, e.g., (The Lazarus group, use, spear-phishing emails).
  4. (4)
    CTI Outline $O=(o_{1},\ldots,o_{n})$: A tactic-centric outline of the TTPs described in the CTI report, where $o_{k}$ is a textual summary of the techniques and procedures associated with $t_{k}$.
AttackSeq-Tactic. This task tests the model’s ability to deduce a tactic $t_{k}\in S$, given a question $Q$ and four shuffled candidate tactics $C=\{c_{r}:r\in[1,4]\}$. The model must select the correct tactic $c_{l}\in C$ that corresponds to $t_{k}$.
AttackSeq-Technique. This task evaluates the model’s capability to infer a technique $e_{j,k}\in E(t_{k})$, given a question $Q$ and four shuffled candidate techniques $C=\{c_{r}:r\in[1,4]\}$. The model must select the correct technique $c_{l}\in C$ that corresponds to $e_{j,k}$.
AttackSeq-Procedure. This task challenges the model to determine the plausibility of a group of procedures $P(e_{j,k})\subseteq P$. The input is a question $Q$ and two candidate choices $C=\{c_{yes},c_{no}\}$. The model must select the correct choice $c_{l}\in C$ that indicates if $P(e_{j,k})$ is plausible. We further divide the task based on the correct answer to the question, i.e. AttackSeq-Procedure-Yes and AttackSeq-Procedure-No.
Within AttackSeqBench, we create three distinct benchmark settings (refer to Section 4.1) that have varying inputs for each benchmarking task differ to the LLMs. These benchmark settings are as follows: Zero-shot, where the input is a question $Q$ and candidate choices $C$; RAG-empowered, where the input is a question $Q$ and candidate choices $C$, along with a list of relevant ATT&CK TTPs to $Q$; Regular, where the input is a question $Q$ and candidate choices $C$, along with a masked CTI outline where the textual summary of $t_{k}$ is removed from $O$, i.e., $O_{masked}=O\setminus o_{k}$.
Figure 2. The pipeline of our dataset construction process. Given the CTI reports, we first leverage a LLM to extract the attack sequence and re-write the CTI outlines. Then we generate massive Q&A pairs with a question generator. Finally, we design the tailored evaluation criteria to enable LLM to select the appropriate Q&A pairs and construct the datasets for AttackSeqBench.
3.2. Dataset Construction
As illustrated in Figure 2, the automated dataset construction pipeline is divided into three phases: (1) Attack Sequence Construction, (2) Question Generation, and (3) Self-Refinement.
3.2.1. Attack Sequence Construction
AttacKG+ (Zhang et al., 2025b) has summarized 500 real-world CTI reports from multiple intelligence sources into tactic-centric outlines, and extracted the associated TTPs by leveraging a LLM-based framework. To construct attack sequences that resemble real-world adversarial behavior, we exclude CTI outlines which contains less than two ATT&CK tactics as they are unlikely to detail attack patterns observed in cyber attacks (Zhang et al., 2025b). Afterwards, we combine the CTI outlines and the associated TTPs from 408 filtered CTI reports to form attack sequences.
3.2.2. Question Generation
We leverage the question generation capabilities of LLMs (Alam et al., 2024; Zhang et al., 2024b; Mucciaccia et al., 2025) to formulate the questions in our QA dataset based on the attack sequences constructed. More specifically, we propose an answer-aware question generation approach, by utilizing GPT-4o (OpenAI, 2024b) to generate a question based on each TTP present with respect to the scope of each benchmarking task for a given attack sequence $S$. To generate more relevant questions, we leverage LLMs’ in-context learning (ICL) ability (Dong et al., 2024) by providing the associated outline $O$ as contextual knowledge, and include few-shot examples based on the example procedures in the ATT&CK KB (refer to Section C.1). For the MCQ tasks, we adopt a rule-based approach to randomly select three distractors from the attack sequence and ATT&CK KB. Due to the limited tactics present in the ATT&CK KB, we randomly sample an adjacent tactic present in the attack sequence i.e., $t_{k+1}$ or $t_{k-1}$, along with two randomly selected tactics from the ATT&CK KB. For AttackSeq-Technique, we select the distractors with reference to the STARC annotation framework (Berzak et al., 2020). The three distractors are as follow: (1) Similar to AttackSeq-Tactic, the first distractor is associated with the same tactic but not supported by the CTI report, i.e., $e_{i,k}\in E(t_{k})$, indicating a plausible misunderstanding of the correct technique $e_{j,k}$; (2) The second distractor refers to a technique that is supported by the CTI report but belongs to another tactic $t_{j}\neq t_{k}$; (3) The third distractor refers to a technique that is not supported by the CTI report and belongs to a randomly chosen tactic.
For the boolean questions, AttackSeq-Procedure, we first generate the questions with correct answer ”Yes”, i.e., for each group of procedures $P(e_{j,k})\subseteq P$. To create the negative questions in AttackSeq-Procedure-No, we randomly sample 70% of the AttackSeq-Procedure-Yes questions from our final QA dataset and prompt the LLM to generate two types of negative questions based on each sampled question (refer to Table 15). Inspired by existing work (Rajpurkar et al., 2018), the two types of negative questions are: (1) We negate the temporal prepositions, i.e., changing ”before” to ”only after” and/or ”after” to ”only before”, such that the sequence of adversarial behavior is incorrect; (2) We replace $P(e_{j,k})$ in the question’s body with a group of procedures associated with a randomly sampled technique that is not supported by CTI report.
3.2.3. Self-Refinement
Inspired by the Self-Refine framework (Madaan et al., 2023), we adopt an iterative three-step approach using GPT-4o to improve the quality of the generated questions. We first assess the quality of the question based on the Answerability aspect (Fu et al., 2024), filtering out the questions that do not satisfy this criteria. More specifically, we check if there is direct evidence in the CTI outline that supports the correct answer, while clearly standing out as the best answer choice. Within this aspect, we also check if the correct answer can be inferred even if the associated summary to the correct answer’s tactic is removed from the outline (Huang et al., 2019), i.e., using $O_{masked}$ to answer the question. Based on the filtered answerable questions, we prompt the LLM to consider the quality of the question and its answer choices based on the following five aspects:
  • Clarity. We check if the question precise and unambiguous. More importantly, we also ensure that question avoid directly mentioning the correct answer such that the inference is required.
  • Logical. We check if the sequence described in the question follow the order of tactics present in the attack sequence.
  • Relevance. We check if the TTPs described in the question directly relate to the attack sequence.
  • Consistency. We check if the question is consistent with the associated TTP that is used for question generation.
  • Answer Consistency. We check if the question can be fully answered by the correct answer, without any contradictions and inconsistencies.
Based on these aspects, we prompt the LLM to give a numerical score out of 5 and select the questions that have a full score of 25. This filter-feedback-refine process is repeated once more to ensure better quality of questions. Refer to Appendix C.2 for the detailed prompts.
Table 1. Number of questions and proportional decrease in each benchmarking task before and after Self-Refinement.
Task Num. (Before) Num. (After) % Decrease
AttackSeq-Tactic 2,158 1,697 -21.36
AttackSeq-Technique 2,937 1,917 -34.73
AttackSeq-Procedure 4,642 2,635 -43.24
3.2.4. Dataset Statistics
The number of generated questions for each benchmark task, along with the corresponding percentage decrease, is shown in Table 1. More details of the dataset, such as the distribution of ATT&CK tactics and techniques associated with the questions, is in Appendix A.1.
3.3. Dataset Evaluation
3.3.1. Human Evaluation
We ask two postgraduate students majoring in computer science, who are knowledgeable about cybersecurity and the MITRE ATT&CK framework, to assess the quality of questions in AttackSeqBench. We randomly sample 35 questions from each of the three benchmarking tasks, including the AttackSeq-Procedure-No subtask, resulting in a total of 140 questions for manual evaluation. Each annotator evaluates the six aspects described in Section 3.2.3 using a 5-point Likert scale, where higher scores indicate better alignment (detailed definitions of each aspect are in Appendix A.2). Based on Table 8, the average accuracy score of annotators over the three tasks is 0.625, suggesting that questions in AttackSeqBench are notably challenging for domain experts. Furthermore, the human evaluation results indicate that the generated questions closely satisfy with each aspect of the evaluation criteria, with an average score of 4.620/5 across the six aspects. In particular, the Logical and Relevance aspects suggests that LLMs are able to construct questions that closely align with attack sequences described in CTI reports.
3.3.2. Automatic Evaluation
To alleviate the laborious task of manual human annotation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge frameworks in aligning with human preferences within domain-specific fields, including cybersecurity (Xu et al., 2024b). We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated questions. Similar to Section 3.3.1, we carefully design prompts for each of the six aspects using a 5-point Likert scale (refer to Appendix C.3 for the detailed prompt). More specifically, we instruct GPT-4o to refer to the evaluation criterion’s definition, along with description of the correct answer from the ATT&CK KB, and rate the question out of 5 for each aspect. Based on Table 9, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences. While the automatic evaluation scores are lower than the human evaluation scores, this implies that it is more challenging for LLMs to answer the questions correctly as compared to domain experts.
4. Experiment
Figure 3. The illustration of three benchmark settings. (a): Regular setting, (b): Zero-shot setting, and (c): RAG-empowered setting.
4.1. Benchmark Settings
As illustrated in Figure 3, we elaborate on the three proposed benchmark settings, Regular, Zero-Shot and RAG-Empowered. The detailed prompts for each setting is in Appendix C.4.
Regular Setting. Inspired by existing works (Ma et al., 2023; Jin et al., 2024), we propose this benchmark setting to assess the impact of contextual information in analyzing attack sequences. We adopt an ICL-based approach by carefully designing a well-constructed prompt using CoT. Within the prompt, we first instruct the LLM to analyze the attack sequence described in the masked CTI outline associated with the question, and subsequently deduce the correct answer based on the given attack sequence.
Zero-Shot Setting. Aside from excelling in semantic understanding, previous works have also demonstrated LLMs’ impressive zero-shot reasoning ability across various downstream tasks (Hou et al., 2024; Kojima et al., 2022). To evaluate the LLMs’ capabilities in zero-shot attack sequence analysis, we omit the masked CTI outline from the prompt in the Regular setting. Instead, we instruct the LLM to select the most plausible answer choice based on the attack sequence described in the question’s body.
RAG-Empowered Setting. Retrieval Augmented Generation (RAG) has gained traction due to its ability to enhance LLMs’ reliability and mitigate hallucinations (Zhang et al., 2025a). We introduce this benchmark setting to assess the impact of providing the relevant ATT&CK TTPs in the question’s body as contextual information, by leveraging a retrieval-based approach and the ATT&CK KB. We first crawl the description and the example procedures of each TTP from the Enterprise ATT&CK Matrix  [^1] https://attack.mitre.org/matrices/enterprise/. Next, we split and embed the text chunks alongside their associated ATT&CK tactic and technique as metadata. Afterwards, we retrieve relevant chunks based on the question’s body and extract the associated ATT&CK TTPs from the metadata. Lastly, we instruct the LLM to utilize the retrieved TTPs and deduce the correct answer. By doing so, we enable the LLM to focus on determining the correct attack sequence rather than interpreting the TTPs in the question.
Table 2. Performance (Accuracy) comparison of various open-source and proprietary LLMs (both fast-thinking and slow-thinking reasoning LLMs) under three settings (i.e., Regular, Zero-Shot, RAG) on three datasets (i.e., Tactic, Technique, Procedure). For each column, bold number indicates the best performance, and underlined number corresponds to the second best performance.
LLMs AttackSeq-Tactic AttackSeq-Technique AttackSeq-Procedure
Regular Zero-Shot RAG Regular Zero-Shot RAG Regular Zero-Shot RAG
Fast-thinking LLMs
[…]
Qwen-2.5-7B-Instruct 0.5121 0.4903 0.4761 0.6693 0.6568 0.6067 0.6584 0.5184 0.4941
Llama-3.1-8B-Instruct 0.4744 0.5085 0.4926 0.6260 0.6333 0.5827 0.6577 0.5328 0.5230
ChatGLM-4-9B-Chat 0.4885 0.4979 0.5009 0.6275 0.6109 0.6030 0.641 0.5408 0.5131
Llama-3.3-70B-Instruct 0.6588 0.5551 0.5681 0.7058 0.6797 0.7037 0.6903 0.5469 0.5279
Qwen-2.5-72B-Instruct 0.5793 0.5863 0.5657 0.5430 0.7162 0.6959 0.7188 0.6285 0.6030
GPT-4o-mini 0.6517 0.6005 0.5692 0.7387 0.7058 0.7021 0.6968 0.5491 0.5340
GPT-4o 0.6093 0.5740 0.5787 0.6755 0.6995 0.7188 0.7359 0.6755 0.6353
Slow-thinking Reasoning LLMs
[…]
DeepSeek-R1-Distill-Qwen-32B 0.5421 0.5698 0.5504 0.5879 0.5816 0.5816 0.6945 0.6258 0.5852
QWQ-32B-Preview 0.5345 0.3377 0.4638 0.5112 0.3918 0.5342 0.7036 0.5696 0.5457
GPT-o3-mini 0.4643 0.5445 0.5215 0.5373 0.5915 0.5822 0.6854 0.6877 0.6459
4.2. Experimental Setting
4.2.1. Implementation Details
To demonstrate the robustness of AttackSeqBench, we evaluate eight fast-thinking LLMs: Mistral-7B-Instruct-v0.3, Qwen-2.5-7B-Instruct, Llama-3.1-8B-Instruct, ChatGLM-4-9B-Chat, Llama-3.3-70B-Instruct, Qwen-2.5-72B-Instruct, GPT-4o-mini and GPT-4o. Additionally, we assess four slow-thinking reasoning LLMs: DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-32B, QWQ-32B-Preview and GPT-o3-mini. For the RAG-empowered benchmark setting, we utilize a hybrid retriever due to its considerable performance in existing work (Zhang et al., 2024b). We combine Okapi BM25 (Robertson and Zaragoza, 2009) as the sparse retriever and a dense retriever based on BGE-EN-ICL (Li et al., 2024) (shortened as BGE) and Chroma DB  [^2] https://www.trychroma.com/, along with a re-ranker. The detailed implementation is in Appendix B.1.
4.2.2. Evaluation Metrics
We evaluate the Accuracy (abbreviated as Acc in the following sections) for all three tasks in AttackSeqBench (Liu et al., 2024; Zhang et al., 2024b; Alam et al., 2024). Specifically, given the total number of questions $M$ and the number of questions $n$ with the correct answer choice chosen by the LLM, the accuracy can be calculated as $Acc=n/M$.
4.3. Performance Comparison
We conduct comprehensive analysis using eight fast-thinking LLMs and four slow-thinking reasoning LLMs in Table 2 and elaborate on three key findings: (1) None of the LLMs clearly outperforms other LLMs across all the tasks. While GPT-4o demonstrate strong performance in AttackSeq-Technique and AttackSeq-Procedure, its performance falls behind Llama-3.3-70B-Instruct and GPT-4o-mini in AttackSeq-Tactic. This implies the complexity and the diversity of our three sub-tasks to some extent. (2) Among these three tasks, we observe that LLMs collectively demonstrate the lowest overall performance for the AttackSeq-Tactic task. This implies LLMs struggle to determine the high-level objectives in an attack sequence, even with the CTI outline given in the context. This may arise due to the abstract nature of the descriptions of tactics which causes the LLM to misinterpret the correct tactic. (3) The AttackSeq-Procedure task yields the highest performance results across all the LLMs. This suggests that LLMs are able to leverage its procedural knowledge to deduce the plausibility of a group of procedures in an attack sequence. However, this may also be attributed by the LLMs’ tendency to favor ”yes” in boolean questions when given insufficient context (Wen et al., 2024), which causes high accuracy for AttackSeq-Procedure-Yes. We explore this phenomenon further in Section 4.4.1.
4.4. Model Analysis
Table 3. Performance comparison of four representative LLMs between AttackSeq-Procedure-Yes and AttackSeq-Procedure-No sub-tasks in each benchmark setting. For each column, the bold value indicates the best performance.
LLMs AttackSeq-Procedure-Yes AttackSeq-Procedure-No
Regular Zero-Shot RAG Regular Zero-Shot RAG
Llama-3.1-8B 0.7899 0.9518 0.9460 0.5432 0.1700 0.1565
GPT-4o 0.7146 0.9403 0.9542 0.7542 0.4462 0.3591
R1 (Llama-8B) 0.8332 0.9526 0.9477 0.4341 0.1161 0.1062
GPT-o3-mini 0.5789 0.7760 0.8242 0.7776 0.6112 0.4915
4.4.1. Impact of Contextual Information
To comprehensively investigate the impact of the contextual knowledge in the given CTI outlines for analyzing attack sequences in LLMs, we compare the performance of LLMs between the Regular and Zero-Shot benchmark settings. Based on in Table 2, we observe that both fast-thinking and slow-thinking reasoning LLM’s accuracy generally increases across each task in the  Regular setting. Among the fast-thinking LLMs, the performance of models with smaller number of parameters, e.g., Llama-3.1-8B-Instruct (shortened as Llama-3.1-8B), have minimal or no improvement for AttackSeq-Tactic and Attack-Technique tasks. This indicates that the baseline LLMs with smaller parameters may not fully leverage the capabilities of ICL within these benchmarking tasks. Similarly for slow-thinking reasoning models, DeepSeek-R1-Distill-Llama-8B (shortened as R1 (Llama-8B)) has no improvement in the Regular setting as compared to the Zero-Shot setting. On the other hand, Llama-3.3-70B-Instruct achieves comparable performance against the closed source models, i.e., GPT-4o and GPT-4o-mini. Among the three tasks, AttackSeq-Procedure has the highest improvement in LLMs from the Zero-Shot setting to the Regular setting. Based on Table 3, we observe a significant increase in the AttackSeq-Procedure-No sub-task. This suggests that the contextual information found in the masked CTI outline is essential in justifying procedures that are unlikely to happen.
Table 4. Performance comparison between three embedding models. Eeb. M. refers to the embedding model used, while Tac, Tec and Pro refers to the three tasks in AttackSeqBench (i.e., Tactic, Technique, and Procedures) respectively. For each row, the bold value indicates the best performance.
#Eeb. M. Task Llama-3.1-8B GPT-4o R1 (Llama-8B) GPT o3-mini
BGE Tac 0.4926 0.5787 0.4532 0.5215
Tec 0.5827 0.7188 0.5138 0.5822
Pro 0.5230 0.6353 0.4968 0.6459
OpenAI Tac 0.4915 0.4814 0.4632 0.5333
Tec 0.5790 0.7178 0.5039 0.5962
Pro 0.5195 0.6330 0.4979 0.6353
ATT&CK- BERT Tac 0.4985 0.5622 0.4590 0.5292
Tec 0.5790 0.7100 0.5018 0.5874
Pro 0.5245 0.6209 0.4953 0.6398
4.4.2. Effectiveness of RAG Strategies
We analyze the capabilities of the RAG-empowered setting based on the results of AttackSeqBench in Table 2. Within the set of fast-thinking LLMs, GPT-4o achieves exceptional performance when leveraging RAG in AttackSeq-Tactic and AttackSeq-Technique tasks. Moreover, the LLMs exhibit the best performance in the AttackSeq-Technique among the three tasks in Table 2. This indicates that LLMs are able to leverage the retrieved TTPs to determine the correct technique. However, the RAG-empowered setting does not consistently yield effective results, as models that have less than or equal to 9B parameters have no improvement as compared to the Zero-Shot setting. This suggests that the information retrieved, i.e., the semantically similar ATT&CK TTPs based on the question’s body, is insufficient for smaller models. Furthermore, we evaluate the performance of four representative LLMs using two additional embedding models in Table 4. The two models are: OpenAI’s text-embedding-3-large (OpenAI, 2024c); ATT&CK BERT (Abdeen et al., 2023) which is an open-source embedding model fine-tuned for cybersecurity. Based on Table 4, BGE acheives the best performance across the three benchmarking tasks.
Figure 4. Performance and model size comparison of nine representative LLMs in the AttackSeq-Tactic task. The size of bubble corresponds to inference time, where horizontal lines denote fast-thinking LLMs and vertical lines indicate slow-thinking reasoning LLMs.
4.4.3. Computational Complexity Analysis
As illustrated in Figure 4, we extensively compare the performance, model size and inference cost of nine open-source LLMs in the AttackSeq-Tactic task. We have the following observations: (1) In general, these LLMs exhibit the performance trend in sequential attack pattern understanding which aligns with the scaling law (Kaplan et al., 2020). (2) Comparing the performance of LLMs with different thinking types, we observe that slow-thinking LLMs cannot significantly and consistently outperform the other. One possible reason is that given the question and masked CTI outlines, fast-thinking LLMs can already deduce the correct answer without any elaborate reasoning. (3) Regarding the inference cost, we notice that slow-thinking LLMs do not incur more reasoning costs compared to fast-thinking LLMs. This prompts an intuitive insight: the small-scale slow-thinking reasoning LLM fine-tuned using AttackSeqBench may enhance its capability in attack sequence analysis significantly while maintaining the lower reasoning cost.
4.4.4. Sampling Parameters Analysis
We assess the impact of two sampling parameters on the performance of LLMs in AttackSeq-Tactic under the Regular setting, namely temperature and maximum output tokens as illustrated in Figure 5. We observe that increasing the temperature at different three intervals has minimal improvements for Llama-3.3-70B-Instruct, while resulting in negative performance for the remaining three LLMs. This indicates that the most optimal performance occurs when the temperature is kept at 0.0. On the other hand, we observe a considerable improvement when increasing the number of tokens for slow-thinking reasoning models. The accuracy of Deepseek-R1-Distill-Llama-8B improved by 8.46% as the maximum output tokens increased from 1024 to 4096, while Deepseek-R1-Distill-Qwen-32B achieve a 24.9% improvement for the same increment. However, further increasing the token limit to 8192 yields diminishing returns for both models. This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in slow-thinking models while considering the associated costs incurred (Wang et al., 2024).
Figure 5. Parameter analysis on temperature (a) and maximum output tokens (b) in AttackSeq-Tactic task using the Regular setting, we perform analysis on four representative LLMs using different colors to represent them.
Figure 6. The performance comparison of four representative LLMs on each tactic in each benchmarking task.
Table 5. Reasoning path comparison between slow-thinking reasoning and fast-thinking LLMs in Zero-Shot setting based on the AttackSeq-Procedure task in Zero-Shot setting.
Question: After delivering spearphishing emails with spyware disguised as a PDF file to gain initial access, is it likely the authors of Hacking Team’s spyware established communication with the identified C&C servers before using VMProtect for obfuscation to evade detection? Answer Choices: A) Yes  B) No (Correct Answer)
DeepSeek-R1-Distill-Qwen-32B Qwen-2.5-72B-Instruct
First, spearphishing emails are emails that look legitimate, which are actually malicious. In this case, the PDF is a disguise for spyware. So the first step is the delivery of the malicious payload via email. (Knowledge Retrieval) Once the victim opens the PDF, the spyware is executed. The next step in an attack chain is usually establishing communication with a command and control (C&C) server. Now, the question mentions VMProtect being used for obfuscation. However, sometimes obfuscation can be applied after deployment to maintain persistence or evade detection during operation. (Breakdown of Attack Sequence) But in the context of the attack sequence described, the spyware has already been delivered and executed. The next logical step would be for the spyware to reach out to the C&C server to report back and receive instructions. (Self-Correction) If the attackers were to use VMProtect at this point, it might be to re-obfuscate the spyware or to protect it further, but that seems a bit out of order. (Reasoning using Attack Sequence) So the answer would be No. (Correct Conclusion) The sequence includes: Delivering spearphishing emails with spyware disguised as a PDF file to gain initial access. Establishing communication with C&C (Command and Control) servers. Using VMProtect for obfuscation to evade detection. (Breakdown of Attack Sequence) The spearphishing emails with a PDF file are used to gain initial access to the target system. (T1566.002: Spearphishing Attachment). Once the initial access is gained, the malware typically establishes communication with the C&C servers to receive further instructions or exfiltrate data. (T1071: Application Layer Protocol). VMProtect is a tool used to obfuscate the malware to evade detection by security software. (T1027: Obfuscated Files or Information). (Knowledge Extraction) The logical sequence of events would be: 1. **Initial Access:**, 2. **C&C Communication:**, 3. **Obfuscation:**. (Incorrect Attack Sequence Order) Based on the logical sequence of TTPs, it is likely that the authors of Hacking Team’s spyware established communication with the C&C servers before using VMProtect for obfuscation to evade detection. Final Answer: Yes (Incorrect Conclusion)
4.5. In-depth Analysis
4.5.1. Impact of TTP Temporal Position
To better understand the relationship between LLMs and attack sequences, we conduct fine-grained analysis on the performance of four representative LLMs under the Regular benchmark setting and across the ATT&CK tactics (refer to Figure 6). Additionally, we identify four categories to group the tactics into overarching attack phases as shown in Figure 6: Initial Intrusion, Exploitation, Stealth Expansion, Objective Orchestration. Our findings are: (1) Based on the standard deviation range across various tactics in each LLM, we observe that the range is the largest for AttackSeq-Tactic, indicating that the LLM’s performance largely varies for this task across different tactics. In contrast, the SD range is the smallest for AttackSeq-Procedure which suggests more stability in the LLMs’ performance. This highlights the challenge for LLMs to determine the correct tactics in attack sequences, as compared to procedural knowledge needed to correctly determine techniques and procedures. (2) GPT-4o consistently outperforms other models across most tasks, particularly in the in the Exploitation and Stealth Expansion phases. On the other hand, Llama-3.1-8B-Instruct and GPT o3-mini achieve comparable performance against GPT-4o in Initial Intrusion and Objective Orchestration phases, especially at the start and end of the sequence (i.e., Reconnaissance and Impact respectively). This inconsistency suggests that TTPs within certain ATT&CK tactics are more challenging for the LLM to interpret, likely due to the rarely observed tactics in real-world CTI reports. More detailed distribution of the performance can be found in Section B.2 of the Appendix.
4.5.2. Inference Path Analysis on Slow-thinking Reasoning LLMs
We perform a case study based on a question from AttackSeq-Procedure-No and compare the difference in the reasoning process in the output of two open-source LLMs: DeepSeek-R1-Distill-Qwen-32B and Qwen-2.5-72B-Instruct in the Zero-Shot setting. We choose not to use GPT-o3-mini in our analysis as the Batch API does not return the reasoning tokens. Based on Table 5, we observe that both models are able to breakdown the attack sequence and extract the relevant attack patterns. The reasoning output tokens suggest that DeepSeek-R1-Distill-Qwen-32B is able to retrieve relevant knowledge based on the attack sequence described in the question. However, Qwen-2.5-72B-Instruct outputs the incorrect order of attack patterns as stated in the question. Furthermore, it highlights the slow-thinking reasoning LLMs’ ability to perform Self-Correction (Kumar et al., 2024), enabling the model to elicit stronger reasoning capabilities and justify why the attack sequence stated in the question is incorrect. Notably, the slow-thinking reasoning LLM considers both incorrect and correct attack sequences before deciding on a conclusion, highlighting the potential in slow-thinking reasoning LLMs in zero-shot tasks.
5. Outlook and Conclusion
Outlook. AttackSeqBench extends beyond existing CRU-related benchmarks by emphasizing sequential patterns in cyber attacks, opening up numerous research avenues for related researchers. For instance, AttackSeqBench can be seamlessly utilized for the complex attack path modeling task as we provided in Section 4.5.1. Moreover, AttackSeqBench can also be leveraged to recognize the attack intention due to its temporal characteristic. Inspired by the ubiquitous causality in real-world scenarios, we believe that integrating the logical attack sequence with the causal inference is a promising research direction, discovering the critical causal relationships within the attack sequences and enabling more accurate predictions of the attackers’ potential action. Given the predicted attack intentions and next action, applying our AttackSeqBench to optimize the defense strategy is not only intuitive but also reasonable. Finally, the constructed dataset can be reformulated as the instruction-tuning data, facilitating the post-training phase of future security-related LLMs.
Conclusion. In this work, we introduced AttackSeqBench, a novel benchmark designed to assess the LLMs’ ability to analyze attack sequences from real-world CTI reports. We propose an automated dataset construction pipeline, enabling our benchmark to be easily extended as cyber threats continue to evolve. Our benchmark incorporates three tasks that encompasses the varying aspects of adversarial behavior, allowing a comprehensive analysis of the LLMs’ understanding and reasoning capabilities in attack sequence analysis. We performed extensive experiments on a diverse set of fast-thinking and slow-thinking reasoning LLMs, offering crucial insights into the abilities and limitations of LLMs in understanding attack sequences. Our work opens up a meaningful new direction towards LLM-driven CRU in real-world cybersecurity operations.
Appendix A Dataset
A.1. Dataset Distribution
The distribution of questions in AttackSeqBench is shown in Table 6, we construct an almost equal proportions of questions in each task to allow fair evaluation in our benchmark.
Based on Figure 7, we observe that the top three most frequent tactics occur in the middle of attack sequences, while the bottom two tactics occurs at the start and end of the attack sequence. The most frequent technique belongs to the Command and Control tactic, this is likely due to the common operations of APTs. Similarly, the most frequent ATT&CK technique is T1071-Application Layer Protocol  [^3] https://attack.mitre.org/techniques/T1071/, which is associated with the Command and Control tactic.
Figure 7. Visualization of the distribution of top 10 tactics and techniques in our AttackSeqBench.
A.2. Dataset Evaluation
Inspired by existing work (Fu et al., 2024), we evaluate the quality of our QA dataset based on six dimensions below:
  • Answerability: We check if there is direct evidence in the CTI outline that supports the correct answer, while clearly standing out as the best answer choice. Within this aspect, we also check if the correct answer can be inferred even if the associated summary to the correct answer’s tactic is removed from the CTI outline.
  • Clarity. We check if the question precise and unambiguous. More importantly, we also ensure that question avoid directly mentioning the correct answer such that the inference is required.
  • Logical. We check if the sequence described in the question follow the order of tactics present in the attack sequence.
  • Relevance. We check if the TTPs described in the question directly relate to the attack sequence.
  • Consistency. We check if the question is consistent with the associated TTP that is used for question generation.
  • Answer Consistency. We check if the question can be fully answered by the correct answer, without any contradictions and inconsistencies.
We give detailed instructions to the two human evaluators by defining each score in the 5-point Likert scale for each aspect (refer to Table 7). The same Likert scale definitions are used during automatic evaluations in the prompt. The detailed results of human evaluation and automatic evaluation are shown in Table 6 and Table 7, respectively. While the automatic evaluation results are lower than human evaluation, the human evaluation shows that most QA pairs in the dataset satisfy the requirements of all aspects. This suggests that automatic evaluation is still limited in knowledge-intensive domains such as in cybersecurity. Note that for AttackSeq-Procedure-No, we only assess the questions in the following four aspects: Answerability; Clarity, Consistency; Answer Consistency.
A.3. Ethics Statement
This study adheres to ethical research practices by focusing on publicly available CTI reports and ensuring that no proprietary or sensitive information is used in the benchmark construction. The automated dataset generation process is designed to maintain the integrity and accuracy of adversarial behavior sequences without fabricating or misrepresenting cyber threats. Furthermore, human evaluation is conducted with careful consideration of annotator expertise and potential biases, ensuring fairness and reliability in assessment. Our work aims to contribute to the cybersecurity community by advancing LLM capabilities while upholding ethical standards in data usage, evaluation, and responsible AI development.
Appendix B Experiments
B.1. Baseline LLMs
To demonstrate the robustness of our proposed datasets, we evaluate eight fast-thinking LLMs and four slow-thinking reasoning LLMs on our benchmark tasks. We leverage vLLM (Kwon et al., 2023) to run all open-source LLMs locally, utilizing two Nvidia H100 GPUs. For the proprietary LLMs, we utilize OpenAI’s Batch API  [^4] https://platform.openai.com/docs/guides/batch to reduce inference cost and perform inference in batches. In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1. Below is the list of LLMs used in our experiments:
Fast-thinking LLMs:
  • Mistral-7B-Instruct-v0.3 Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) is an instruct fine-tuned version of the Mistral-7B-v0.3 which exhibits extended vocabulary and supports the v3 tokenizer.
  • Qwen2.5-7B-Instruct Qwen2.5-7B-Instruct (Team, 2024a) is one of the advanced Qwen series LLMs. It exhibits better instruction following and long text generating capability.
  • LLaMa-3.1-8B-Instruct LLaMa-3.1-8B-Instruct (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
  • LLaMa-3.2-3B-Instruct LLaMa-3.2-3B-Instruct (Grattafiori et al., 2024) is a lightweight 3B version of LLaMA-3.2, prioritizing the rapid response and the deployment on low-resource devices.
  • Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct (Team, 2024a) is one of the latest LLMs from Qwen and serves as the biggest Qwen-2.5 series LLM with 72B parameters.
  • LLaMa-3.3-70B-Instruct LLaMa-3.3-70B-Instruct (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF).
  • GPT-4o GPT-4o (OpenAI, 2024b) is one of the most advanced closed-source LLMs, which is a multi-lingual and multi-modal language model developed and functions well in real-time processing  [^5] To assess its superiority, we leverage the most recent GPT-4o-2024-11-20 and GPT-4o-mini-2024-07-18 (at the time point when we did this work) as the benchmark evaluator, respectively..
  • GPT-4o-mini gpt-4o-mini (OpenAI, 2024a) is a compact variant of the GPT-4o (OpenAI, 2024b), which is optimized for faster inference and cost efficiency 5.
Slow-thinking reasoning LLMs:
  • DeepSeek-R1-Distill-Llama-8B DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025) is fine-tuned from the open-source Llama-3.1-8B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
  • DeepSeek-R1-Distill-Qwen-32B DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025) denotes the distilled version from the well-trained Qwen2.5-32B without any Reinforce Learning stage.
  • QWQ-32B-Preview QWQ-32B-Preview (Team, 2024b) is a preview release which gives LLM time to ponder, to question, and to reflect, enabling the deeper insight into complex problems.
  • GPT-o3-mini GPT-o3-mini (OpenAI, 2025) is designed with a focus on enhancing LLMs’ reasoning capabilities. It leverages the Chain of Thought (CoT) to break down complex problems into several simpler steps to achieve this objective.
The retrievers in our experiments are built using LlamaIndex (Liu, 2022). We set the number of retrieved chunks $u$ and its size $l$ to be $3$ and $512$ respectively.
B.2. TTP Temporal Position Analysis
We conduct a detailed breakdown of the performance in four indicative LLMs based on each ATT&CK tactic associated to the questions in AttackSeqBench. Based on Table 10, the mean and standard deviation (SD) for each tactic and between each benchmark task is shown. We observe that the overall highest standard deviation occurs within the AttackSeq-Tactic task for all models, indicating that it is challenging for LLMs to correctly answer questions in AttackSeq-Tactic.
Table 10. Results of performance of four representative models on all the 14 tactics, and the associative statistics of mean and standard deviation (SD).
Tactic
Reconna-
issance
Resource
Devel.
Initial
Access
Execu. Persist.
Privilege
Escala.
Defense
Evasion
Credent.
Access
Discov.
Lateral
Move.
Collec.
Cmd. &
Control
Exfiltr. Impact Mean SD
Tactic Llama-3.1-8B 0.6170 0.3077 0.4406 0.5573 0.4260 0.3023 0.3867 0.4211 0.2551 0.5778 0.4390 0.6444 0.5915 0.4194 0.4561 0.1239
GPT-4o 0.6383 0.5000 0.4368 0.7470 0.5207 0.4884 0.5430 0.5000 0.7245 0.7111 0.6951 0.7155 0.7042 0.6774 0.6144 0.1095
R1 (Llama-8B) 0.4681 0.2692 0.3448 0.5455 0.3609 0.2791 0.2188 0.3421 0.2857 0.4667 0.3415 0.7238 0.5070 0.3548 0.3934 0.1350
GPT-o3-mini 0.5319 0.4615 0.4269 0.5635 0.3018 0.1395 0.3711 0.3289 0.4490 0.5111 0.3537 0.7071 0.493 0.6774 0.4512 0.1499
Mean 0.5638 0.3846 0.4123 0.6033 0.4024 0.3023 0.3799 0.3980 0.4286 0.5667 0.4573 0.6977 0.5739 0.5323 - -
SD 0.0786 0.1132 0.0454 0.0961 0.0938 0.1434 0.1325 0.0792 0.2149 0.1066 0.1643 0.0362 0.0971 0.1697 - -
Technique Llama-3.1-8B 0.6556 0.4878 0.5327 0.5890 0.7324 0.4667 0.5914 0.5625 0.6839 0.5179 0.6557 0.6884 0.7500 0.7241 0.6170 0.0943
GPT-4o 0.6222 0.6341 0.6262 0.6568 0.7254 0.6444 0.6512 0.6000 0.7513 0.6429 0.6148 0.7329 0.7632 0.6897 0.6682 0.0541
R1 (Llama-8B) 0.5667 0.4634 0.4393 0.5551 0.5775 0.4444 0.4651 0.5250 0.6321 0.3929 0.5246 0.6027 0.7368 0.4828 0.5292 0.0911
GPT-o3-mini 0.5222 0.5366 0.4836 0.5085 0.6241 0.4889 0.4333 0.3875 0.5389 0.6250 0.5328 0.6473 0.7237 0.6552 0.5505 0.0933
Mean 0.5917 0.5305 0.5205 0.5774 0.6649 0.5111 0.5353 0.5188 0.6516 0.5447 0.5820 0.6678 0.7434 0.6380 - -
SD 0.0591 0.0755 0.0802 0.0624 0.0764 0.0907 0.1031 0.0927 0.0896 0.1153 0.0638 0.0557 0.0170 0.1072 - -
Procedure Llama-3.1-8B 0.6634 0.5319 0.6489 0.6552 0.6667 0.5556 0.6809 0.6331 0.6987 0.6633 0.6027 0.6606 0.6906 0.6140 0.6404 0.0489
GPT-4o 0.7030 0.6170 0.8351 0.7586 0.7469 0.6852 0.7325 0.7554 0.7067 0.7041 0.6712 0.7515 0.777 0.7193 0.7260 0.0523
R1 (Llama-8B) 0.6535 0.5957 0.6064 0.6681 0.6728 0.6111 0.6474 0.5540 0.6027 0.5918 0.5023 0.6485 0.6547 0.5439 0.6109 0.0510
GPT-o3-mini 0.7030 0.7234 0.8085 0.6853 0.6852 0.7037 0.6991 0.6835 0.6640 0.6633 0.6256 0.6707 0.7338 0.5439 0.6852 0.0585
Mean 0.6807 0.6170 0.7247 0.6918 0.6929 0.6389 0.6900 0.6565 0.6680 0.6556 0.6005 0.6828 0.7140 0.6053 - -
SD 0.0260 0.0796 0.1139 0.0462 0.0368 0.0684 0.0355 0.0848 0.0473 0.0467 0.0714 0.0467 0.0530 0.0829 - -
B.3. Error Analysis
To better understand the limitations of LLMs in analyzing attack sequences, we examine a specific cases where GPT-4o successfully and fails to answer the given question correctly (refer to Table 11). In the first example, the model incorrectly associates the technique in answer choice A, i.e., T1574-Hijack Execution Flow, with the Execution tactic. While this technique involves the execution of malicious payloads based on the ATT&CK KB, it is associated to only the following tactics, Persistence, Privilege Escalation and Defense Evasion. This incorrect interpretation suggests that the model struggles in distinguishing the inherent ambiguity found in TTP descriptions, which affects their ability to analyze attack sequences.
Table 11. Error analysis of GPT-4o in AttackSeq-Technique. The LLM incorrectly associates technique T1574 with Execution tactic
Question & Answer Choices LLM Response
Question: Before establishing persistence through creating a Run key in the registry, which ATT&CK technique most likely facilitated the execution of the backdoor FinalBot.exe after the WMI discovery process identified the presence of antivirus software? Answer Choices: A) T1574-Hijack Execution Flow B) T1003.001-LSASS Memory C) T1059.008-Network Device CLI D) T1129-Shared Modules (Correct Answer) **A: T1574-Hijack Execution Flow** This technique involves hijacking the execution flow of a legitimate process to execute malicious code. DLL sideloading is a common method for this technique. The CTI report explicitly mentions that Sidewinder used DLL sideloading (‘cmpbk32.dll‘ to execute ‘Srvstr.exe‘), which aligns with this technique. This likely facilitated the execution of the backdoor ‘FinalBot.exe‘ after the WMI discovery process. **B: T1003.001-LSASS Memory**: This technique involves dumping credentials from the LSASS process memory. There is no mention in the CTI report of credential dumping or any activity related to LSASS. This choice is irrelevant. **C: T1059.008-Network Device CLI**: This technique involves using a command-line interface (CLI) to interact with network devices. The CTI report does not mention any interaction with network devices or the use of a CLI. This choice is irrelevant. **D: T1129-Shared Modules** This technique involves executing malicious code by loading shared modules. While DLL sideloading involves shared modules, the specific ATT&CK technique for DLL sideloading is T1574-Hijack Execution Flow. This choice is less precise than option A. The most plausible sequence of TTPs involves the use of **T1574-Hijack Execution Flow** (via DLL sideloading) to execute the backdoor ‘FinalBot.exe‘ after the WMI discovery process and before establishing persistence (Wrong Conclusion).
Appendix C Prompts Strategy
C.1. Question Generation Prompts
We utilize few-shot prompting for each of our benchmark tasks, the prompts for AttackSeq-Tactic, AttackSeq-Technique, AttackSeq-Procedure-Yes, AttackSeq-Procedure-No is shown in Tables 12, 13, 14, 15 respectively.
Table 12. Few-shot prompt for question generation for AttackSeq-Tactic. We insturct the LLM to generate a question based on the given CTI outline, where the answer to the question is the given ATT&CK TTP.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, where the answer to the question is the given MITRE ATT&CK tactic. The question should focus on inferring the given tactic by using the attack sequence based on the remaining tactics in the CTI report. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is the given MITRE ATT&CK tactic. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, which ATT&CK tactic most likely occurred before Ke3chang achieved persistence by adding a Run key? Answer: Execution - Question: Which ATT&CK tactic most likely occurred before Axiom gained initial access to the victim’s network using SQL injection? Answer: Resource Development - Question: Which ATT&CK tactic most likely occurred after Ke3chang establishes connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: Exfiltration - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, which ATT&CK tactic most likely occurred before APT41 used Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: Execution [End of Examples]
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic given below.
Table 13. Few-shot prompt for question generation for AttackSeq-Technique.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. 3. A **MITRE ATT&CK Technique** present in the CTI report. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, where the answer to the question is the given MITRE ATT&CK technique that belongs to the given ATT&CK tactic. The question should focus on inferring the given technique by using the attack sequence based on the remaining tactics in the CTI report. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is the given MITRE ATT&CK technique. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, which ATT&CK technique most likely occurred before Ke3chang achieved persistence by adding a Run key? Answer: T1059-Command and Scripting Interpreter - Question: Which ATT&CK technique most likely occurred before Axiom gained initial access to the victim’s network using SQL injection? Answer: T1583.002-DNS Server - Question: Which ATT&CK technique most likely occurred after Ke3chang establishes connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: T1020-Automated Exfiltration - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, which ATT&CK technique most likely occurred before APT41 used Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: T1569.002-Service Execution [End of Examples]
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic and technique given below.
Table 14. Few-shot prompt for question generation for AttackSeq-Procedure-Yes.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive four parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. 3. A **MITRE ATT&CK Technique** present in the CTI report. 4. 4. A list of **Procedures** present in the CTI report, where each procedure is represented as a (Subject, Relation, Object) triplet. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, the question should focus on inferring the given list of procedures based on the given MITRE ATT&CK tactic and technique. The answer to the question is ”Yes”, indicating that the given list of procedures is likely to occur in the attack sequence. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is ”Yes”. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, is it likely that the Ke3chang malware will run commands on the command-line interface before achieving persistence by adding a Run key? Answer: Yes - Question: Is it likely that Axiom will acquire dynamic DNS services for use in the targeting of intended victims before gaining initial access to the victim’s network using SQL injection? Answer: Yes - Question: Is Ke3chang likely to perform frequent and scheduled data exfiltration from compromised networks after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: Yes - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, is APT41 likely to use Windows services to execute DUSTPAN before using Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: Yes [End of Examples]
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic, technique and procedures given below.
Table 15. Few-shot prompt for question generation for AttackSeq-Procedure-No.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **Reference Question-Answer Pair** that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always ”Yes”. 2. A **Reference MITRE TTP** that is NOT supported in the CTI report. [End of Inputs]
Your task is to generate two questions based on the given reference question, such that attack sequence described in the question is modified and the correct answer to the two questions is ”No”. The definitions of the two questions are as follows: 1. Question 1 should negate the ”before” and/or ”after” clauses of the reference question, such that the attack sequence contradicts the original sequence of TTPs in the reference question. 2. Question 2 should replace the main procedures in the reference question with the provided reference MITRE TTP, such that the replaced main procedures are not found in the CTI report. [Steps]: Please follow these steps: 1. Analyze the Reference Question-Answer Pair: - Identify and outline the attack sequence in the order presented in the reference Question-Answer Pair. 2. Construct the Questions: - Design two questions that modify the attack sequence described in the reference question. - Ensure that the answer to the questions is ”No”. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the examples below for guidance. [Examples]: Example Questions: - Example 1: Reference Question: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely run commands on the command-line interface before achieving persistence by adding a Run key? Reference Answer: Yes Reference TTP: Tactic: Initial Access, Technique: T1651-Cloud Administration Command, Example Procedures: AADInternals can execute commands on Azure virtual machines using the VM agent. APT29 has used Azure Run Command and Azure Admin-on-Behalf-of (AOBO) to execute code on virtual machines. Pacu can run commands on EC2 instances using AWS Systems Manager Run Command. Question 1: After achieving persistence by adding a Run key, will the Ke3chang malware run commands on the command-line interface only after gaining initial access through compromised VPN accounts? Answer: No Question 2: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely execute commands on Azure virtual machines using the VM agent before achieving persistence by adding a Run key? Answer: No - Example 2: Reference Question: Will Axiom acquire dynamic DNS services for use in the targeting of intended victims before gaining initial access to the victim’s network using SQL injection? Reference Answer: Yes Reference TTP: Tactic: Resource Development, Technique: T1585.001-Social Media Accounts, Example Procedures: APT32 has set up Facebook pages in tandem with fake websites. Cleaver has created fake LinkedIn profiles that included profile photos, details, and connections. EXOTIC LILY has established social media profiles to mimic employees of targeted companies. Question 1: Will Axiom acquire dynamic DNS services for use in the targeting of intended victims only after gaining initial access to the victim’s network using SQL injection? Reference Answer: No Question 2: Will Axiom set up Facebook pages in tandem with fake websites before gaining initial access to the victim’s network using SQL injection? Answer: No - Example 3: Reference Question: Will Ke3chang perform frequent and scheduled data exfiltration from compromised networks after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Reference Answer: Yes Reference TTP: Tactic: Exfiltration, Technique: T1030-Data Transfer Size Limits, Example Procedures: AppleSeed has divided files if the size is 0x1000000 bytes or more. APT28 has split archived exfiltration files into chunks smaller than 1MB. APT41 transfers post-exploitation files dividing the payload into fixed-size chunks to evade detection. Question 1: Will Ke3chang perform frequent and scheduled data exfiltration from compromised networks only before establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: No Question 2: Will Ke3chang divide files if the size is 0x1000000 bytes or more after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: No [End of Examples]
3. Provide the Question-Answer Pairs: - Please follow the output format: ”Question 1: ¡insert question 1 here¿ Answer: ¡insert answer to question 1 here¿ Question 2: ¡insert question 2 here¿ Answer: ¡insert answer to question 2 here¿”.
[End of Steps] Following the steps above, please generate two questions based on the Reference Question-Answer Pair and Reference MITRE TTP given below. Please only provide the final output of the two questions.
C.2. Dataset Refinement Prompts
The prompt templates to filter based on the Answerability criteria is in Table 16, while the feedback and refinement prompts are in Table 17 and Table 18 respectively.
Table 16. Prompt template for verifying Answerability during Self-Refinement.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **TTP Description**: A reference description of the correct answer corresponding to the question. 3. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. [End of Inputs] Your task is to evaluate the answerability of the given question using the provided information in the CTI Outline and TTP description. We define answerability based on three factors below: 1. The correct answer must be supported by the CTI outline. 2. The correct answer must clearly stand out as the best answer choice to the question based on the CTI outline. 3. Suppose the masked_tactic paragraph is removed from the CTI outline, the correct answer must be deducible from the answer choices by using the information provided in remaining tactics of the CTI outline and TTP description. You may also refer to your external cybersecurity knowledge to determine if the correct answer is deducible. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Match the correct answer with the provided TTP description. - Determine step-by-step if the question is answerable based on the definition above. 3. Output evaluation result: - Output one of the following: - ”A”: Indicates that the question is answerable. - ”B”: Indicates that the question is not answerable. - ”C”: Indicates that you do not know/cannot determine if the question is answerable. - Please also include a short and concise explanation of your evaluation result. - Please follow the output format: ”Explanation: ¡insert explanation here¿ Evaluation Result: ¡insert letter here¿”. [End of Steps] Following the steps above, please evaluate the question using the CTI report and description below and only output the evaluation result.
Table 17. Prompt template for feedback on question quality based on evaluation criteria.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the Tactics, Techniques and Procedures (TTPs) in MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **TTP Description**: A reference description of the correct answer to the question. 3. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI outline, along with one correct answer and distractors among the answer choices. [End of Inputs]
[Evaluation Criteria]: Your task is to evaluate the QA pair and provide your feedback for each of the criteria defined below: 1. **Clarity**: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. **Logical**: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. **Relevance**: Does the question directly relate to the CTI outline? 4. **Consistency**: Does the question align with the provided TTP Description? 5. **Answer Consistency**: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI outline: - Read the CTI outline carefully. - Identify and outline the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Assess each criterion step by step, rating it on a scale of 1 to 5 (1 = poor, 5 = excellent). - Provide a short and concise feedback for each rating. 4. Output Feedback Scores: - Please follow the output format: Feedback Scores: - Clarity: ¡Your feedback¿ (¡Score¿/5) - Logical: ¡Your feedback¿ (¡Score¿/5) - Relevance: ¡Your feedback¿ (¡Score¿/5) - Consistency: ¡Your feedback¿ (¡Score¿/5) - Answer Consistency: ¡Your feedback¿ (¡Score¿/5) Total Score: ¡Total Score¿/25 [End of Steps] Following the steps above, please evaluate the Question with Answer Choices below using the provided CTI report and TTP Description. Please only output the Feedback Scores.
Table 18. Prompt template for question refinement.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. **Feedback Results**: A list of feedback scores and explanations for each desired criterion of the question defined below. [End of Inputs] [Evaluation Criteria]: Please refer to the definition of each feedback criterion: 1. **Clarity**: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. **Logical**: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. **Relevance**: Does the question directly relate to the CTI outline? 4. **Consistency**: Does the question align with the provided TTP Description? 5. **Answer Consistency**: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? Your task is to iteratively refine the quality of the given question based on the feedback provided in the Feedback Results. [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. 2. Analyze the Question with Answer Choices: - Read the question and the provided answer choices carefully. 3. Analyze the Feedback Results: - Based on the feedback given in each criterion, refine the question to improve the each aspect. - Please ensure that the correct answer to the refined question is the same as the original question. - Please also ensure that the question avoids hinting at the correct answer. 4. Output the Refined Question: - Please follow the output format: ”Refined Question: ¡Your refined question here¿”. [End of Steps] Following the steps above, please refine the question based on the Feedback Results and CTI Outline provided below. Please only output the refined question.
C.3. Automatic Evaluation Prompt
We utilize the definitions in the evaluation criteria in Table 7 to create prompts. We show an example prompt template for evaluating the Logical aspect of the question shown in Table 19.
Table 19. Prompt template for evaluating the Logical aspect.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. **Description to Correct Answer**: The description of the correct answer from the MITRE ATT&CK framework. [End of Inputs] Your task is to rate the question on one metric below. [Definition of Logical Aspect] Evaluation Criteria: Logical (1-5): Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding and/or subsequent tactics in the CTI outline such that it logically leads to the correct answer? The scale is defined as follows: 1 - Not Logical: The question does not align with the logical sequence of MITRE ATT&CK tactics in the CTI outline. It ignores or contradicts the natural order of tactics and TTPs. 2 - Weak Logical Alignment: The question shows minimal alignment with the MITRE ATT&CK sequence. It may reference unrelated tactics or disrupt the logical flow. 3 - Moderately Logical: The question has some logical alignment, but it may not reference preceding or subsequent tactics clearly. The sequence could be improved. 4 - Strong Logical Alignment: The question follows the expected sequence of MITRE ATT&CK tactics and references preceding or subsequent TTPs in a logical manner. 5 - Perfect Logical Alignment: The question perfectly aligns with the MITRE ATT&CK framework, referencing relevant TTPs in a way that naturally leads to the correct answer. [Steps:] Evaluation Steps: 1. Analyze the CTI report and Description to Correct Answer: - Read the report and the provided description carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Evaluate the Question: - Read the question and the provided answer choices carefully. - Using the CTI outline and provided description to the correct answer, Rate the question on a scale of 1-5 according to the evaluation criteria above. 3. Output evaluation score: - Please only output the numerical evaluation score based on the defined criteria. [End of Steps] Following the steps above, please evaluate the question and only output the numerical evaluation score.
C.4. Answering Prompts
The prompt templates for the three benchmarking tasks, Regular, Zero-Shot and RAG-empowered are shown in Tables 20, 21, 22 respectively.
Table 20. Prompt template for Regular setting.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **Question** about a sequence of TTPs with several answer choices. [End of Inputs]
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the CTI report. Note that the CTI report contains key details required for your analysis, but it may not directly state the answer. Your evaluation of the answer choices is essential to arrive at the correct answer. [Steps:] Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the Question: - Read the question and its answer choices. - Identify the sequence of TTPs mentioned in the question. 3. Compare and Evaluate: - Match the extracted attack sequence from the CTI report with the details in the question. - Evaluate each answer choice to determine which one aligns best with the attack sequence and any critical contextual information. 4. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below using the provided CTI report.
Table 21. Prompt template for Zero-Shot setting.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive a question about a sequence of TTPs with several answer choices. Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question. [End of Inputs]
[Steps:] Please follow these steps: 1. Analyze the Question: - Read the question and its answer choices. - Identify the sequence of TTPs mentioned in the question. 2. Compare and Evaluate: - Evaluate each answer choice to determine which one aligns best with the attack sequence in the question. 3. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below.
Table 22. Prompt template for RAG-empowered setting.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **Question** about a sequence of TTPs with several answer choices. 2. A list of **Related TTPs** that are relevant to the question. [End of Inputs]
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question.
[Steps:] Please follow these steps: 1. Analyze the Question: - Carefully read the question and its answer choices. 2. Analyze the Related TTPs: - Analyze the list of Related TTPs to understand the context of the question. 3. Compare and Evaluate: - Based on the related TTPs, evaluate each answer choice to determine which one aligns best with the attack sequence in the question. 4. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below.